Jo Ann and her lab partner are completing a free fall lab in physics. Her partner drops the ball in front of the motion sensor to measure its speed. Assuming acceleration is constant and air resistance is negligible, how far does the object fall if it reaches a speed of 9.8 m/s?

1
Expert's answer
2022-05-01T17:32:28-0400

1.

"s=\dfrac{gt^2}{2}, v=gt"

Then

"s=\dfrac{v^2}{2g}=\dfrac{(9.8m/s)^2}{2(9.8m/s^2)}=4.9m"